OpenStudy (anonymous):

evaluate (125/27)^2/3

OpenStudy (anonymous):

i like to start off by taking the \[\sqrt[3]{125/27}\]

OpenStudy (anonymous):

then i break it down into \[\sqrt[3]{125}/\sqrt[3]{27}\]

OpenStudy (anonymous):

which equals what?

OpenStudy (anonymous):

i honestly dont know how to evaluate, thanks for the help though.

OpenStudy (anonymous):

okay the cubic root means what value times itself three times = 125...for instance x*x*x=8 what would x be? 2 bc 2*2*2=8 so x*x*x=125? what is x

OpenStudy (anonymous):

Oh! okay, i understand now, thank you so much!

OpenStudy (anonymous):

after you get that dont forget to take the cubed root of 27 and then square the whole thing

OpenStudy (anonymous):

okay! thank you, again!